OpenStudy (anonymous):

Select the counterexample that makes this conjecture false: For any real number x, x2 ≥ x.

OpenStudy (anonymous):

x = 1/2 x = -2 x = |3| x = 0

OpenStudy (anonymous):

Try plugging in each of those into the equation. And see which one the inequality isn't true for.

OpenStudy (anonymous):

can u help? i forget how to solve inequalitys

OpenStudy (anonymous):

Well, you don't need to do much with the inequality other than realize what it is saying which is: given any x, x^2 is always "Greater Than or Equal To" x. So, by testing each choice, i.e. squaring each one, you will find that only one of them gets smaller when squared.

OpenStudy (anonymous):

There is nothing to solve; it's plug-and-chug.

OpenStudy (anonymous):

The statement is basically saying, "x squared is always bigger than x." You prove this is false by finding a counterexample. In this case, a counterexample is a value for x which when squared, is smaller than just x. Just start squaring things, you'll find the answer!

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

Very welcome!

OpenStudy (anonymous):

i get x = 1/2 :) is that right?